Using Binomial Theorem, give the closed form expression for: "\textstyle\sum_{k=0}^n""{n \choose k}"3n · 2k

1
Expert's answer
2021-11-04T19:58:52-0400

Using Binomial Theorem, let us give the closed form expression for

"\sum\limits_{k=0}^n {n \choose k}3^n · 2^k."


It follows that

"\sum\limits_{k=0}^n {n \choose k}3^n · 2^k
=3^n\sum\limits_{k=0}^n {n \choose k}2^k1^{n-k}\\=3^n(2+1)^n=3^n3^n=3^{2n}=9^n."


We conclude that the closed form expression is "9^n."